# Population Protocols Revisited: Parity and Beyond

### v2

@@ Abstract
The population protocol model is used to study the power of pairwise interactions between simple, indistinguishable entities, referred to as agents. In this model, each agent is equipped with limited storage, represented by a single state drawn from a predefined state space. In the probabilistic variant of population protocols adopted here, a random scheduler selects an interacting pair of agents uniformly at random at each step of the protocol. In this variant, in addition to efficient state utilisation, one is also concerned with the stabilisation time, determined by the number of interactions required to reach the final configuration, divided by the population size $n$. We say that a protocol is time-efficient if it stabilises in polylogarithmic (in $n$) time, and space-efficient if it utilises polylogarithmic state space. We also say that a protocol is stable if it stabilises with the correct answer with probability 1, and silent if agents do not change their states after stabilisation.
For nearly two decades, population protocols have been extensively studied, yielding efficient solutions for central problems in distributed computing, including leader election, and majority computation, a predicate type in Presburger Arithmetic closely tied to population protocols [5, 7]. Surprisingly, no protocols have achieved both time- and space-efficiency for congruency predicates, including parity computation, which are complementary in this arithmetic framework. This gap highlights a significant challenge in the field [16, 19, 21, 24]. To address this gap, we explore the parity problem, where agents are tasked with computing the parity of the given sub-population size. Then we extend the solution for parity to compute congruences modulo an arbitrary $m$.
@@
Previous research on efficient population protocols has focused on protocols that minimise both stabilisation time and state utilisation for specific problems. In contrast, this work slightly relaxes this expectation, permitting protocols to place less emphasis on full optimisation and more on universality, robustness, and probabilistic guarantees. This allows us to propose a novel computing paradigm that integrates population weights (or simply weights), a robust clocking mechanism, and efficient anomaly detection coupled with a switching mechanism (which ensures slow but always correct solutions). This paradigm facilitates universal design of efficient multistage stable population protocols. Specifically, the first efficient parity and congruence protocols introduced here use both $O(\log^{3}n)$ states and achieve silent stabilisation in $O(\log^{3}n)$ time. We conclude by discussing the impact of implicit conversion between unary and binary representations enabled by the weight system, with applications to other problems, including the computation and representation of (sub-)population sizes, a problem posed in [21].
@@
All algorithmic solutions presented in this paper are stable, silent, and guarantee stabilisation time with high probability (whp) defined as $1-n^{-\eta},$ for a constant $\eta>0.$
@@ 1 Introduction
The model of population protocols originates from the seminal paper [5], providing tools suitable for the formal analysis of pairwise interactions between simple, indistinguishable entities referred to as agents. These agents are equipped with limited storage, communication, and computation capabilities. When two agents engage in a direct interaction, their states change according to the predefined transition function, which forms an integral part of the population protocol. The weakest assumptions in population protocols concern the fixed (constant-size) operational state space of agents, with the population size $n$, whether exact or approximate, not hard-coded into either the states or the transition function. It is assumed that a protocol starts in the predefined initial configuration of agents’ states representing the input, and it stabilises in one of the final configurations of states representing the solution to the considered problem. In the standard probabilistic variant of population protocols adopted here, in each step of a protocol, the random scheduler selects an ordered pair of agents: the initiator and the responder, which are drawn from the whole population uniformly at random. The lack of symmetry in this pair is a powerful source of randomness utilised by population protocols. In the probabilistic variant, in addition to efficient state utilisation, one is also interested in the time complexity, where the sequential time refers to the number of interactions leading to the stabilisation of a protocol in a final configuration. More recently, the focus has shifted to the parallel time, or simply the time, defined as the sequential time divided by the size $n$ of the whole population. The (parallel) time reflects on the parallelism of simultaneous independent interactions of agents utilised in efficient population protocols that stabilise in time $O(\mbox{poly}\log n)$. All algorithmic solutions presented in this paper are stable (always correct), silent (agents do not change their states after stabilisation), and guarantee stabilisation time with high probability (whp) defined as $1-n^{-\eta},$ for a constant $\eta>0.$
In this paper, the rules of a transition function are denoted as $A+B\rightarrow A^{\prime}+B^{\prime}$, meaning that if there is an interaction initiated by any agent in state $A$, with any agent in state $B$, then the former changes its state to $A^{\prime}$, while the latter changes its state to $B^{\prime}$. Some of the protocols presented in this work operate on compound states, which means that an agent state is a concatenation of some sets of substates $S_{1},\ldots,S_{k}$, and defining these sets defines the full state space as $S=S_{1}\times\ldots\times S_{k}$. Encoding a state of an agent can be seen as labeling it by a number represented in a binary form. Encoding compound states can be achieved by utilising specified positions in binary representation for each substate set.
@@ 1.1 Related Work and Results
Population protocol studies primarily follow two main research avenues. The first focuses on understanding the behavior of large populations driven by simple interaction rules. The second treats population protocols as computational systems, investigating their computational power based on factors such as the type of scheduler and the memory available to individual agents. This study refers to the latter, focusing on designing efficient population protocols with robust probabilistic guarantees.
The original population protocol model [5] assumes that the number of states and the transition function are constant, i.e., independent of the population size $n$. However, for fundamental problems such as leader election [23], majority computation [1], and the evaluation of various functions and predicates [9], it has been demonstrated that stable, fixed-state protocols operating in sublinear time are infeasible. This limitation has motivated the study of population protocols in which the number of states is permitted to grow with $n$, enabling such problems to be solved efficiently, i.e., in polylogarithmic time.
Prior work on efficient population protocols has emphasized minimizing both stabilisation time and state usage for specific problems. In contrast, this study adopts a broader approach, prioritizing universality, robustness, and probabilistic guarantees over full optimisation. We introduce a novel computing paradigm that incorporates population weights (or simply weights), a robust clocking mechanism, and efficient anomaly detection paired with a switching mechanism to ensure slow but consistently correct solutions. This new paradigm supports the universal design of efficient multistage protocols, which are stable (always correct), silent (agents don’t alter their states after stabilisation), and ensure stabilisation in time O(polylog n)O($\text${polylog }n) with high probability.
@@
A (one-way) epidemic process [6] is a simple yet effective protocol in which each agent can be in one of two possible states, denoted as 0 (susceptible) or $1$ (infected), with the transition function consisting of a single significant rule $(x)+(y)\rightarrow(x)+(\max\{x,y\})$.
@@
The epidemic process establishes a fundamental $\Theta(\log n)$-time communication bound in population protocols. It lies at the core of numerous algorithmic tools and solutions in population protocols, including phase clocks [2, 6, 8, 22, 26] and, in turn, fast leader election [3, 15, 2, 14, 8, 25, 13], majority computation [4, 1, 2, 15, 29, 30, 11, 10, 22], and others. A recent paper [27] presents a generalisation of the epidemic process to the $k$-contact epidemic model, in which an agent becomes infected only after interacting with $k$ distinct infected agents.
@@
A central challenge in distributed computing is symmetry breaking, typically achieved by electing one agent as the leader. In the context of population protocols, the results in [17, 20] laid the foundation for proving that leader election cannot be solved in sublinear time using a fixed number of states [23]. Alistarh and Gelashvili [3] then introduced an upper bound with a protocol stabilising in $O(\log^{3}n)$ time using $O(\log^{3}n)$ states per agent. Later, Alistarh et al. [1] explored general trade-offs between state and time complexity, establishing a separation between slowly stabilising protocols using $o(\log\log n)$ states and rapidly stabilising ones requiring $O(\log n)$ states. Subsequent advances improved these bounds: Bilke et al. [15] achieved $O(\log^{2}n)$ time whp with $O(\log^{2}n)$ states; Alistarh et al. [2] and Berenbrink et al. [14] reduced states to $O(\log n)$ while preserving $O(\log^{2}n)$ time whp using synthetic coins; Gąsieniec and Stachowiak [8] further lowered state usage to $O(\log\log n)$ with the same time bound whp; Gąsieniec et al. [25] traded high probability for expected $O(\log n\log\log n)$ time; and finally, Berenbrink et al. [13] achieved the optimal expected $O(\log n)$ time.
@@
In this paper, a direct use of leader election via pairwise duels appears in the slow parity protocol, see Section 2.1. In fast protocols, the elected leader initiates weight creation and coordinates rounds and process synchronisation. Our leader election mechanism builds on coin-flip reduction across successive rounds [6, 8, 13], with round numbering ensuring that at least one leader survives. In the rare event of multiple surviving leaders, anomaly detection resolves the conflict in expected $O(n)$ time, see Section 3.3.
Majority Computation
@@
Let $A$ and $B$ be subpopulations of $n$ agents. The majority problem is to determine whether $|A|>|B|$, $|A|<|B|$, or $|A|=|B|$. For efficient majority computation, Angluin et al. [6] introduced an approximate majority protocol: if the initial bias $|A|-|B|=\omega(\sqrt{n}\log n)$, it converges to the majority opinion in $O(\log n)$ time whp. Condon et al. [18] improved the required bias to $\Omega(\sqrt{n\log n})$. For the exact majority, Alistarh et al. [4] gave the first stable polylogarithmic-time protocol using $\Omega(n)$ states. Subsequent work [1, 2, 15, 29, 30, 11, 10] improved the space-time trade-off, culminating in an optimal $O(\log n)$-time protocol with $O(\log n)$ states [22]. This matches the $\Omega(\log n)$ state lower bound of Alistarh et al. [2] for stable sublinear-time protocols under monotonicity and output dominance, both satisfied by [22]. Non-stable protocols (allowing positive error probability) achieve faster convergence whp with less memory. Berenbrink et al. [11] gave a protocol with $O(\log\log n)$ states and $O(\log^{2}n)$ time. Also, Kosowski and Uznański [28] achieved $O(1)$ states and $O(\mathrm{polylog}\,n)$ time whp.
@@
In this paper, we use exact majority computation as a conceptual balance scale, see Figure 2, enabling comparison of sub-population cardinalities through appropriately chosen weight combinations. Like other modern majority protocols, ours also uses cancelling (opposite biased agents $\to$ unbiased) and splitting (biased $\to$ two biased), doubling the bias per phase until $n$. As indicated earlier, our protocols prioritise universality, robustness, and probabilistic guarantees over full optimisation. For details, see Section 3.4.
Figure 1: Balance scale. [^2] This is an AI-generated picture.
@@
In this paper, synchronisation is achieved via round counting by each agent, reinforced by leader-attributed fine-grained interaction counting and epidemic waves. This precise time-stamping system enables protocols with efficient anomaly detection through execution-time inconsistencies, essential for the MC+ protocols.
MC+ Protocols
Our new computing paradigm relies on Monte Carlo protocols enhanced with anomaly detection, termed MC+ protocols. These either silently stabilise to the correct answer or detect efficiently an anomaly, triggering a slow (polynomial-time) but always correct solution. We distinguish between clock anomalies, caused by desynchronisation, and protocol anomalies, which arise from interactions between agents in conflicting states. We show that MC+ protocols are suitable building blocks for constructing more complex solutions. For detail see Section 3.
@@
In the population protocol model, integer values are represented in unary format, where each value corresponds to the cardinality of a subpopulation of agents in a designated state. This representation makes addition straightforward, i.e., fast and stable. While it is well understood that fixed-state Monte Carlo population protocols can simulate a limited variant of register machines (recognising problems in LOGSPACE) [6], the efficient and stable computation of basic operations such as subtraction (multiplication and division) remains significantly more challenging. The primary obstacle in subtraction lies in the need for effective pairing of agents from two subpopulations, which may require linear time when their cardinalities are comparable or equal. A similar challenge arises in majority protocols, where it is addressed by amplifying the difference (bias) between the cardinalities to a magnitude that population protocols can process efficiently [6]. In the same paper, the authors propose a method for efficient Monte Carlo type subtraction by iteratively generating powers of two, combined with simple addition and more sophisticated majority protocols. This approach serves as a prototype for our system of weights, which introduces an implicit binary representation of integer values, supported by corresponding operations.
@@
While this paper primarily focuses on the parity problem, our new computing paradigm also enables efficient computation of congruences, by providing a suitable framework which facilitates universal design of efficient multistage stable population protocols (Section 3.2). Specifically, the first efficient parity (Theorem 1) and congruence (Theorem 9) protocols are proved correct, both using $O(\log^{3}n)$ states and achieving stabilisation in $O(\log^{3}n)$ time (Section 4). Both protocols are stable (always correct), silent (agents do not change their states after stabilisation), and guarantee stabilisation time whp. We conclude (Section 5) by discussing the impact of implicit conversion between unary and binary representations enabled by the weight system, with applications to other problems, including computation and representation of (sub-)population sizes.
To streamline the presentation, we assume that the subpopulations of agents representing input, temporary, and output values are each at most $\frac{n}{6}$. However, our solutions remain correct and efficient even for larger fractions of $n$. We also assume that the upper estimate $LOG$ on $\log n$ is provided with constant-factor accuracy and incorporated into the transition function. Finally, we note that a recent study [19] constructs succinctly encoded slow (polynomial time) population protocols for Presburger arithmetic, including congruences modulo any integer $m$. However, since [19] does not address the silence property, for completeness we include here a natural extension of our parity protocol that achieves silence.

### v3

@@ Abstract
The population protocol model is used to study the power of pairwise interactions between simple, indistinguishable entities, referred to as agents. In this model, each agent is equipped with limited storage, represented by a single state drawn from a predefined state space. In the probabilistic variant of population protocols adopted here, a random scheduler selects an interacting pair of agents uniformly at random at each step of the protocol. In this variant, in addition to efficient state utilisation, one is also concerned with the stabilisation time, determined by the number of interactions required to reach the final configuration, divided by the population size $n$. We call a protocol efficient if it is time‑efficient, stabilising in polylogarithmic time in $n$, and space‑efficient, using a polylogarithmic state space. We also say that a protocol is stable if it stabilises with the correct answer with probability 1, and silent if agents do not change their states after stabilisation.
For nearly two decades, population protocols have been extensively investigated, producing efficient solutions for central problems in distributed computing, notably majority computation, one of the two primary predicates in Presburger arithmetic. Yet the complementary class of congruency predicates has remained substantially more elusive [5, 7]. This long-standing gap constitutes a major open challenge in the area. Namely, efficient (non-stable) solutions exist via Monte-Carlo simulation of register machines [6], but all known stable protocols for these predicates are slow, either exponential time [16, 24] or polynomial [19]. In this paper, we present a new methodology that enables the design of both efficient and stable protocols for computing congruency predicates. As a first step, we focus on the parity problem, in which agents must compute the parity of a designated sub-population size. We then generalise the obtained technique to compute congruence modulo an arbitrary integer $m$.
@@
Our congruency protocols achieve efficiency through a modular construction, with particular emphasis placed on universality, robustness, and strong probabilistic guarantees. Central to this approach is the MC+ computing paradigm, which enables integration of population weights (or simply weights), a robust clocking mechanism, and effective anomaly detection with a switching mechanism that guarantees fallback to a slow yet always-correct stable protocol. This paradigm facilitates universal design of efficient multistage stable population protocols. Specifically, the first efficient parity and congruence protocols introduced here use both $O(\log^{3}n)$ states and achieve silent stabilisation in $O(\log^{3}n)$ time. We conclude by discussing the impact of implicit conversion between unary and binary representations enabled by the weight system, with applications to other problems, including the computation and representation of (sub-)population sizes, a problem posed in [21].
@@
All algorithmic solutions presented in this paper are stable, silent, and guarantee stabilisation time in expectation and with high probability (whp) defined as $1-n^{-\eta},$ for a constant $\eta>0.$
@@ 1 Introduction
The model of population protocols originates from the seminal paper [5], providing tools suitable for the formal analysis of pairwise interactions between simple, indistinguishable entities referred to as agents. These agents are equipped with limited storage, communication, and computation capabilities. When two agents engage in a direct interaction, their states change according to the predefined transition function, which forms an integral part of the population protocol. It is assumed that a protocol starts in the predefined initial configuration of agents’ states representing the input, and it stabilises in one of the final configurations of states representing the solution to the considered problem. In the standard probabilistic variant of population protocols adopted here, in each step of a protocol, the random scheduler selects an ordered pair of agents: the initiator and the responder, which are drawn from the whole population uniformly at random. The lack of symmetry in this pair is a powerful source of randomness utilised by population protocols. In the probabilistic variant, in addition to efficient state utilisation, one is also interested in the time complexity, where the sequential time refers to the number of interactions leading to the stabilisation of a protocol in a final configuration. More recently, the focus has shifted to the parallel time, or simply the time, defined as the sequential time divided by the size $n$ of the whole population. The (parallel) time reflects the parallelism arising from simultaneous independent agent interactions used in efficient population protocols, which stabilise in time $O(\text{poly}\log\,n)$ while also using polylogarithmic state space. In this paper, the rules of a transition function are denoted as $A+B\rightarrow A^{\prime}+B^{\prime}$, meaning that if there is an interaction initiated by any agent in state $A$, with any agent in state $B$, then the former changes its state to $A^{\prime}$, while the latter changes its state to $B^{\prime}$. All algorithmic solutions presented in this paper are stable (always correct), silent (agents do not change their states after stabilisation), and guarantee stabilisation time with high probability (whp) defined as $1-n^{-\eta},$ for an arbitrary fixed constant $\eta>0.$
@@ 1.1 Related Work and Results
Population‑protocol research largely follows two directions: analysing the behaviour of large populations with predefined interaction rules; and studying computational power of population protocols under various schedulers and memory constraints. This work belongs to the latter, focusing on designing efficient protocols with robust probabilistic guarantees. The original population‑protocol model [5] assumes a fixed set of states an transitions, independent of the population size $n$. However, for fundamental tasks such as leader election [23], majority computation [1], two central tools in our computing paradigm, and the evaluation of various functions and predicates [9], it is known that stable, fixed‑state significantly faster than linear protocols are not feasible. This limitation has motivated models in which the number of states may grow with $n$, enabling efficient solutions.
Much of the prior work on efficient stable population protocols has focused on specific problems. This study adopts a broader perspective, emphasising universality, robustness, and probabilistic performance guarantees. We propose the MC+ paradigm, which integrates population weights, a resilient clocking mechanism, and effective anomaly detection paired with a switching strategy that falls back to a slower but always correct protocol when necessary. This paradigm enables the universal construction of efficient multistage protocols that are stable (always correct), silent (agents cease changing state after stabilisation), and achieve stabilisation in time $O(\text{poly}\log\,n)$ whp.
@@
A (one-way) epidemic process [6] is a simple yet effective protocol in which each agent can be in one of two possible states, denoted as 0 (susceptible) or $1$ (infected), with the transition function consisting of a single meaningful rule $(x)+(y)\rightarrow(x)+(\max\{x,y\})$.
@@
The epidemic process establishes a fundamental $\Theta(\log n)$-time communication bound in population protocols. It lies at the core of numerous algorithmic tools and solutions in population protocols, including phase clocks [2, 6, 8, 22, 26] and, in turn, efficient leader election [3, 15, 2, 14, 8, 25, 13], majority computation [4, 1, 2, 15, 29, 30, 11, 10, 22], and others. Recent work [27] introduces the $k$-contact epidemic model, requiring $k$ distinct infected contacts for infection.
@@
A central challenge in distributed computing is symmetry breaking, typically achieved by electing one agent as the leader. In the context of population protocols, the results in [17, 20] laid the foundation for proving that leader election cannot be solved in sublinear time using a fixed number of states [23]. Alistarh and Gelashvili [3] then introduced an upper bound with a protocol stabilising in $O(\log^{3}n)$ time using $O(\log^{3}n)$ states per agent. Later, Alistarh et al. [1] explored general trade-offs between state and time complexity, establishing a separation between slowly stabilising protocols using $o(\log\log n)$ states and rapidly stabilising ones requiring $\Omega(\log\log n)$ states. Subsequent advances improved upper bounds: Bilke et al. [15] achieved $O(\log^{2}n)$ time whp with $O(\log^{2}n)$ states; Alistarh et al. [2] and Berenbrink et al. [14] reduced states to $O(\log n)$ while preserving $O(\log^{2}n)$ time whp using synthetic coins; Gąsieniec and Stachowiak [8] further lowered state usage to $O(\log\log n)$ with the same time bound whp; Gąsieniec et al. [25] traded high probability for expected $O(\log n\log\log n)$ time; and finally, Berenbrink et al. [13] achieved the optimal expected $O(\log n)$ time.
@@
In our efficient protocols, the elected leader initiates weight creation and coordinates rounds and synchronisation. Our leader-election mechanism builds on coin-flip reduction across successive rounds [6, 8, 13], with round numbering ensuring at least one leader survives. In the rare case that multiple leaders remain, anomaly detection resolves the conflict in expected $O(n)$ time (Section 3.3).
Majority Computation and Weights
@@
Let $A$ and $B$ be subpopulations of $n$ agents. The majority problem, a predicate from Presburger Arithmetic, is to determine whether $|A|>|B|$, $|A|<|B|$, or $|A|=|B|$. For efficient majority computation, Angluin et al. [6] introduced an approximate majority protocol: if the initial bias $|A|-|B|=\omega(\sqrt{n}\log n)$, it converges to the majority opinion in $O(\log n)$ time whp. Condon et al. [18] improved the required bias to $\Omega(\sqrt{n\log n})$. For the exact majority, Alistarh et al. [4] gave the first stable polylogarithmic-time protocol using $\Omega(n)$ states. Subsequent work [1, 2, 15, 29, 30, 11, 10] improved the space-time trade-off, culminating in an optimal $O(\log n)$-time protocol with $O(\log n)$ states [22]. This matches the $\Omega(\log n)$ state lower bound of Alistarh et al. [2] for stable sublinear-time protocols under monotonicity and output dominance, both satisfied by [22]. Non-stable protocols (allowing positive error probability) achieve faster convergence whp with less memory. Berenbrink et al. [11] gave a protocol with $O(\log\log n)$ states and $O(\log^{2}n)$ time. Also, Kosowski and Uznański [28] achieved $O(1)$ states and $O(\mathrm{polylog}\,n)$ time whp.
@@
In this paper, we use exact majority computation as a conceptual balance scale based on system of weights formed of disjoint subpopulations $W_{i}$ of agents, where $|W_{i}|=2^{i},$ for all $i,\dots,\log(\frac{n}{6}).$ This enables comparison of sub-population cardinalities through appropriately chosen weight combinations. Like other majority protocols, ours also uses cancelling (opposite biased agents $\to$ unbiased) and splitting (biased $\to$ two biased), doubling the bias per phase until $n$. For details, see Section 3.4.
@@
In this paper, synchronisation is achieved via round counting by each agent, reinforced by leader-attributed fine-grained interaction counting and epidemic waves. This precise time-stamping system enables protocols with effective anomaly detection through execution-time inconsistencies, essential for the MC+ protocols.
MC+ Paradigm
Our new computing paradigm relies on Monte Carlo protocols enhanced with anomaly detection, termed MC+ protocols. These either silently stabilise to the correct answer or detect effectively an anomaly, triggering a slow (polynomial-time) but always correct solution. We distinguish between clock anomalies, caused by desynchronisation, and protocol anomalies, which arise from interactions between agents in conflicting states. We show that MC+ protocols are suitable building blocks for constructing more complex solutions. Because our MC+ protocols for leader election and majority must satisfy additional properties, their complexity does not match the fastest known results. Pursuing optimal bounds here would complicate the presentation and distract from the conceptual contributions. For detail see Section 3.
@@
In the population protocol model, integer values are represented in unary format, where each value corresponds to the cardinality of a subpopulation of agents in a designated state. This representation makes addition straightforward, i.e., efficient and stable. While it is well understood that fixed-state Monte Carlo population protocols can simulate a limited variant of register machines (recognising problems in LOGSPACE) [6], the efficient and stable computation of basic operations such as subtraction (multiplication and division) remains significantly more challenging. The primary obstacle in subtraction lies in the need for effective pairing of agents from two subpopulations, which may require linear time when their cardinalities are comparable or equal. A similar challenge arises in majority protocols, where it is addressed by amplifying the difference (bias) between the cardinalities to a magnitude that population protocols can process efficiently [6]. In the same paper, the authors propose a method for efficient Monte Carlo type subtraction by iteratively generating powers of two, combined with simple addition and more sophisticated majority protocols. This approach serves as a prototype for our system of weights, which introduces an implicit binary representation of integer values, supported by corresponding operations.
@@
While this paper focuses primarily on the parity problem, our MC+ paradigm also enables efficient computation of congruences by supporting the universal design of efficient multistage stable population protocols (Section 3.2).We present the first stable and silent protocols for both parity (Theorem 1) and congruence (Theorem 9), each using $O(\log^{3}n)$ states and stabilising in $O(\log^{3}n)$ time both in expectation and with high probability (Section 4). In Section 5, we discuss the impact of implicit unary-to-binary conversion via the weight system and its applications to other problems, including (sub-)population size computation and representation.
To streamline the presentation, we assume each subpopulation (input, temporary, output) is at most $\frac{n}{6}$; A recent work [19] constructs succinctly encoded, slow (polynomial-time) population protocols for Presburger arithmetic, including modular congruences. Since [19] does not address silence, we include also our generic polynomial congruence protocol.